OpenStudy (anonymous):

A store displays seven computers on a shelf side-by-side. Find the total length of the computers on the table if the first computer was six inches wide and the width of each successive computer is three inches longer than the previous one. Show all your work and describe the sigma notation used in answering the question above

OpenStudy (anonymous):

1st computer=6" each next is 3" longer so 2nd computer=6+3=9" 3rd=9+3=12" 4th=12+3=15" 5th=15+3=18" 6th=18+3=21" 7th=21+3=24" now to find the total length add them up all...as =6+9+12+15+18+21+24=105

OpenStudy (anonymous):

whats the sigma notation

ganeshie8 (ganeshie8):

6, 9, 12,... what type of sequence is this ?

OpenStudy (anonymous):

arthmetic sequence

ganeshie8 (ganeshie8):

yes, can you write the general term ? also called `nth term`

OpenStudy (anonymous):

n=1

OpenStudy (anonymous):

upper index=7

OpenStudy (anonymous):

3n+3(middle)

ganeshie8 (ganeshie8):

first term, \(\large a_1 = 6\) common difference, \(\large d=3\) nth term, \(\large a_n = ?\)

OpenStudy (anonymous):

6+(n-1)3

ganeshie8 (ganeshie8):

\[\large \text{Total length} = \sum \limits_{n=1}^7 a_n\]

ganeshie8 (ganeshie8):

plugin the expression for nth term, \(a_n\) above ^^

OpenStudy (anonymous):

ok thnx

ganeshie8 (ganeshie8):

\[\large \text{Total length} = \sum \limits_{n=1}^7 6+(n-1)*3\]

ganeshie8 (ganeshie8):

simplifying gives you : \[\large \text{Total length} = \sum \limits_{n=1}^7 (3n+3)\]